Joint Versus Separate Returns. Carl has $60,000 of salary and $11,000 of itemized deductions. Carol has $90,000 of salary and $16,200 of itemized deductions. They are married and under age 65. a. Compute their taxable incomes if they file separately and Carol claims itemized deductions on her return. b. Compute their taxable income if they file jointly. Assume their total itemized deductions equals the sum of their separate itemized deductions. Tax Year 2019.
Answer #1

Their taxable incomes if they file separately and Carol claims itemized deductions on her return -

Taxable Income = Salary - Itemised deduction

Carl = $60,000 – $12,000 = $48,000

Carol= $90,000 - $16,200 = $73,800

Their taxable income if they file jointly, assume their total itemized deductions equals the sum of their separate itemized deductions.

Taxable Income = Total salary - Total of itemised deduction

Total Salary = 60000+90000 = $150,000

Total of itemised deducton =12000+16200 = $26,200

Taxable income if filed jointly= $150000-$26200=$123,800